3. The Senior High School researchers claim that more than 20% of Senior High



School male students have tried smoking cigarette. After collecting 150 random



samples, they found that 60 of them have tried smoking cigarette.



Claim:



Parameter



Symbol for parameter



Ho and Ha complementary pair.



Hypotheses in words:



Ho



Ha:



Hypotheses in symbols



Ho



На:



4. In a certain town a school principal hypothesized that students enroll in schools



within 5 km from their homes. To check this claim you ask 38 students from



the said town. You found out that the average distance between the students



home and their schools is 5.6 km.



Claim:



Parameter



Symbol for parameter



Ho and Ha complementary pair.



Hypotheses in words:



Ho



Ha:



Hypotheses in symbols



Ho



На:

1
Expert's answer
2022-05-22T23:50:08-0400

3. The population proportion "p" is the parameter.

Claim "p>0.2."

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:" the population proportion is no most 20%.

"H_a:" the population proportion is greater than 20%.

"H_0:p\le0.2"

"H_a:p>0.2"

This corresponds to a right-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a right-tailed test is "z_c =1.6469."

The rejection region for this right-tailed test is "R = \{z: z>1.6469\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p}{\sqrt{\dfrac{p(1-p)}{n}}}=\dfrac{\dfrac{60}{150}-0.2}{\sqrt{\dfrac{0.2(1-0.2)}{150}}}=6.1237"

Since it is observed that "z =6.1237>1.6469=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>6.1237)=0," and since "p = 0 < 0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p" is greater than 0.2, at the "\alpha = 0.05" significance level.


4.

The average distance between the students home and their schools "\mu" is the parameter.

Claim "\mu=5km."

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:" the average distance is 5 km.

"H_a:" the average distance is not 5 km.

"H_0:\mu=5"

"H_a:\mu\not=5"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.